OpenStudy (anonymous):

One to one function h is defined by h(x)=(3x-4)/(-2x+9). Find h^-1, the inverse of h. Then, give the domain and range of h^-1 using interval notation.

OpenStudy (anonymous):

could have sworn i answered this exact question earlier

OpenStudy (anonymous):

domain of \(h\) is found by setting \(-2x+9=0\) and solving for \(x\)

OpenStudy (anonymous):

Aleks didn't take it satellite. not sure why I know that the inverse is: h^-1=(9x+4)/(2x+3). Just need help finding the domain and range in interval notation.

OpenStudy (anonymous):

maybe it was some sort of syntax error

OpenStudy (anonymous):

you just need the domain and range of \(h^{-1}\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

domain is \[(\infty, -\frac{3}{2})\cup (-\frac{3}{2},\infty)\]

OpenStudy (anonymous):

is that what you put?

OpenStudy (anonymous):

I put (-\[(-\infty, 9/2) union (9/2, \infty)\]

OpenStudy (anonymous):

that is for the range

OpenStudy (anonymous):

man!!! No wonder

OpenStudy (anonymous):

that is the domain of \(h\) and the range of \(h^{-1}\)

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

yw